OpenStudy (anonymous):

1. For the circle with equation (x -4)2 + (y +2)2 = 9, answer each question: a) What are the coordinates of the center? 2 points. b) What are the radius and diameter of the circle? 4 points.

OpenStudy (anonymous):

\[(x-h)^2+(y-k)^2=r^2\] \[(x-4)^2+(y+2)^2=9\] what is \((h,k)\) ?

OpenStudy (anonymous):

As much as I need to learn this I don't have the time right now sadly

OpenStudy (anonymous):

Its timed and I have to submit it and another one by 8

OpenStudy (anonymous):

Well thanks for nothing

OpenStudy (isaiah.feynman):

(h,k) = (4,-2) r=3 D=6